A and B appear for an interview for two posts. The probability of A’s
selection is 1/3 and that of B ‘s selection is 2/5 . Find the probability
that only one of them will be selected.

Grade:12

2 Answers

shubham sharma
10 years ago

7/15.

= P(a)''.P(b) plus P(a).P(b)''. {where P(x)'' =1-P(x)}

jitender lakhanpal
10 years ago

hi ramit

 

there would be 2 cases

1) when A will be selected and not B    OR  2) when B will be selected not A

then applying AND ,OR principle

P(A) . P(Bbar) + P(Abar) P(B)

P(A) = 1/3    P(B) = 2/5

P(Abar) = 2/3     P(Bbar) = 3/5

put them in the equation.

We are  here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

 

Thanks and Regards

Jitender

(askIITians expert)